# The birth mass function of neutron stars

### v1

@@ G Potential selection effects
Here we discuss potential selection effects that might have an impact on the inference of neutron star mass distribution. Our compilation of neutron star mass measurements result from a diverse range of observations, which may introduce different observation biases. For simplicity, we first focus on our major subset of data – masses measured with the radio pulsar timing method. In Extended Data Figure 6, the top panel shows the measured mass versus the observed spin period for 39 recycled pulsars (blue) and other neutron stars (orange). The bottom panel shows the measured mass versus the luminosity at 1.4 GHz calculated from the flux densities and pulsar distances as reported in the ATNF Pulsar Catalogue[95] for 39 recycled pulsars. From both plots, one sees that there is no apparent correlation between the mass and the detectability of a pulsar. This suggests that our inferred neutron star mass distribution is unlikely to be biased by selection effects associated with pulsar surveys.
@@
Another potential selection bias is related to the fact that all but one neutron stars in our compilation are in binary systems [^4] The one isolated neutron star (PSR J0030+0451) is a fully recycled pulsar which was likely to be part of a disrupted binary system.. Since massive stars, the progenitors of neutron stars, are commonly formed in binary systems, the neutron star mass function determined in this work might be a good representation of the overall neutron star population, except that binary systems containing neutron stars seen today are only a fraction of binary systems that have given birth to neutron stars. The supernova explosion associated with the formation of a neutron star might disrupt the binary system. If the survival probability of binary systems is correlated with the neutron star mass, e.g., more massive neutron stars receive larger kicks at birth, then the population of observed binary pulsar systems would be biased toward a particular range of neutron star masses. The effect of such a correlation between the neutron star mass and supernova kicks has been tentatively investigated in population synthesis calculations[96, 97]. In particular, using neutron star birth masses derived from modern, parameterized supernova models, ref.[96] find a neutron star birth mass distribution for the primary in double neutron star binaries that is skewed toward lower masses than the single-star mass function. Conversely, the mass distribution of secondaries is shifted upwards slightly compared to the single-star case (see their figure 7). Whereas the aggregated distribution of neutron star birth masses, in particular its shape, is similar to the single-star case, systematic differences in the slope of the mass distribution in double neutron stars, neutron star-white dwarf binaries and isolated neutron stars cannot be excluded; see, e.g., figure 8a in ref.[97]. This will be an interesting area for future research with a much larger catalog of neutron star mass measurements.
@@
To summarize, while we believe that our results are robust against potential selection effects, we do expect to unveil a fuller picture of neutron star mass function in the next decade or so when we have a sizable sample of mass measurements for every subpopulation of neutron star systems, e.g., binary neutron star mergers from gravitational-wave observations, neutron star-white dwarf binaries from pulsar observations, and isolated neutron stars from microlensing observations[98, 99].
Extended Data Figure 10: Top: the measured mass versus spin period for 39 recycled pulsars (blue) and other neutron stars (orange). Bottom: the measured mass versus luminosity at 1.4 GHz for 39 recycled pulsars, where the horizontal error bars account for uncertainties in the flux density and distance. We adopt the distance estimates from the ATNF Pulsar Catalogue [^6] https://www.atnf.csiro.au/research/pulsar/psrcat/ while assuming a 20% error. In both panels, the error bars are 1-$\sigma$ credible interval, while the measurement uncertainty of spin periods in the upper panel is too small to be seen.

### v2

@@ G Potential selection effects
Here we discuss potential selection effects that might have had an impact on the inference of neutron star mass distribution. Our compilation of neutron-star mass measurements was the result of a diverse range of observations, which may have introduced different observation biases. For simplicity, we first focus on our main subset of data: masses measured with the radio pulsar timing method. Extended Data Figure 6a shows the measured mass versus the observed spin period for 39 recycled pulsars (blue) and other neutron stars (orange). Extended Data Figure 6b shows the measured mass versus the luminosity at 1.4 GHz calculated from the flux densities and pulsar distances as reported in the ATNF Pulsar Catalogue[95] for 39 recycled pulsars. From both plots, one can see that there is no apparent correlation between the mass and the detectability of a pulsar. This indicates that our inferred neutron-star mass distribution is unlikely to be biased by selection effects associated with pulsar surveys.
@@
Another potential selection bias may have arisen because all but one of neutron stars in our compilation are in binary systems [^4] The one isolated neutron star (PSR J0030+0451) is a fully recycled pulsar. It was probably part of a disrupted binary system.. As massive stars, the progenitors of neutron stars, are commonly formed in binary systems, the neutron star mass function determined in this work may be a good representation of the overall neutron star population, except that binary systems containing neutron stars seen today are only a fraction of binary systems that have given birth to neutron stars. The supernova explosion associated with the formation of a neutron star could disrupt the binary system. If the survival probability of binary systems is correlated with the neutron star mass, for example, more massive neutron stars receive larger kicks at birth, then the population of observed binary pulsar systems would be biased towards a particular range of neutron star masses. The effect of such a correlation between the neutron star mass and the supernova kick has been tentatively investigated in population synthesis calculations[96, 97]. In particular, using neutron star birth masses derived from modern, parameterized supernova models, ref.[96] found a neutron-star birth-mass distribution for the primary in double neutron star binaries that is skewed towards lower masses than the single-star mass function. Conversely, the mass distribution of secondaries is shifted upwards slightly compared to the single-star case (see their figure 7). Although the aggregated distribution of neutron star birth masses, in particular its shape, is like that of the single-star case, systematic differences in the slope of the mass distribution in double neutron stars, neutron star-white dwarf binaries and isolated neutron stars cannot be excluded; see, for example, figure 8a in ref.[97]. This will be an interesting area for future research with a much larger catalogue of neutron star mass measurements.
@@
To summarize, although we believe that our results are robust against potential selection effects, we do expect to unveil a fuller picture of neutron-star mass function in the next decade or so when we have a sizeable sample of mass measurements for every subpopulation of neutron star systems, for example, binary neutron star mergers from gravitational-wave observations, neutron star and white dwarf binaries from pulsar observations, and isolated neutron stars from microlensing observations[98, 99].
[…]